OpenStudy (anonymous):

Solve the equation. 4x^2 - 20 = 5

OpenStudy (anonymous):

\[4x^2=20+5=25\] and \[x^2=\frac{ 25 }{ 4 }\] lastly, \[x=\sqrt{\frac{ 25 }{ 4 } }= \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)